OpenStudy (calculusxy):

Simplify to lowest terms: 1/2(2x + y)(4z)

OpenStudy (calculusxy):

@shamil98

OpenStudy (shamil98):

Distribute it out..

OpenStudy (calculusxy):

I pretty much understand how the answer explains it. But I don't get why they wouldn't cross out 2 even if the denominator was 2, but crossed out 4 to 2, making it 2z(2x+y)

OpenStudy (shamil98):

\[\frac{ 1 }{ 2 } (2x+y)(4z) = \frac{ (2x+y)(4z) }{ 2 } = 2z(2z+y)\]

OpenStudy (calculusxy):

Why are there two z's?

OpenStudy (shamil98):

a typo that should be 2z(2x+y)

OpenStudy (calculusxy):

But why shouldn't 2 cross out 2?

OpenStudy (shamil98):

You can take two out of the numerator, so you can cross off the two in the denominator.

OpenStudy (shamil98):

Because there the denominator is reduced to 1.. you haven't anything to reduce it by ..

OpenStudy (shamil98):

\[\frac{ 1 }{ \cancel 2 } (2x+y) (\cancel 4 2z) =(2x+y)(2z)\]

OpenStudy (calculusxy):

Okay. It's clear to me now. Thanks.